OpenStudy (anonymous):

Fourier serie for f(x)=3x^2-2x in -pi

hartnn (hartnn):

i hope u know the formulas ?

OpenStudy (anonymous):

\[a0+\sum_{1}^{\infty}_{n}cosnx+_{n}sinnx\]

hartnn (hartnn):

wow... anyways, could u find a0 =... ?

OpenStudy (anonymous):

an and bn exist before cosx and sinx

OpenStudy (anonymous):

yea with oyler formula!

hartnn (hartnn):

\(\large a_0=(1/\pi)\int \limits_{-\pi}^{\pi}(3x^2-2x)dx=....?\)

OpenStudy (anonymous):

euler formula! bn and an are determined with euler formula too.

OpenStudy (anonymous):

∫−ππ(3x2)d-∫−ππ(2x)d=...? yea?

hartnn (hartnn):

i don't know how we use euler here, i would integrate by parts for an and bn

hartnn (hartnn):

u go on,and tell what u finally get for a0=....?

hartnn (hartnn):

yes, its 1/2 pi i thought he wrote a0/2 in the formula, which i generally do.

OpenStudy (anonymous):

for find the a0 , u should separate ∫(3x2−2x)d to two ∫(3x2)dx-∫2x)dx.

hartnn (hartnn):

yeah, that was correct...what u got a0 finally ?

OpenStudy (anonymous):

a0 = 2(pi)^2

hartnn (hartnn):

\(\large a_0=(1/2\pi)\int \limits_{-\pi}^{\pi}(3x^2-2x)dx=(1/2\pi)[x^3-x^2]^{\pi}_{-\pi}\)

OpenStudy (anonymous):

3x^2 is an even function. and other one is Individual function.

OpenStudy (anonymous):

we should use this state i think!

OpenStudy (anonymous):

@hartnn what about other parts?!

hartnn (hartnn):

to find an and bn, you can use integration by parts. try it...if u get stuck, i'll help.

OpenStudy (abb0t):

Some good questions being asked today :)

OpenStudy (anonymous):

@abb0t good solution is need for these lol

hartnn (hartnn):

why don't u try by yourself ?

hartnn (hartnn):

we'll correct u if u get it wrong..

OpenStudy (anonymous):

i tried on paper, it is so long to type here. :D

hartnn (hartnn):

final step u got ?

OpenStudy (anonymous):

i separated that integration to two integration and find a0 an and bn for each of them. so sum them. and after that put it in initial formula!

hartnn (hartnn):

what u got a0=...?an=...?bn=...?

OpenStudy (anonymous):

an for first integration is this: an=\[\int\limits_{0}^{\pi}3x^{2}cosnxdx \rightarrow......?\]

hartnn (hartnn):

did u miss 1/pi in the beginning ?

OpenStudy (anonymous):

2/pi i missed in beginning! sry

hartnn (hartnn):

so u having trouble with integration by parts ?

OpenStudy (anonymous):

Fractional integrals is needed to solve this?

hartnn (hartnn):

fractional integers ? no...

OpenStudy (anonymous):

so how can u find the that integration?!

hartnn (hartnn):

didn't u try integration by parts ? u know what that is, right ? \(\int uv = ... ?\)

hartnn (hartnn):

here u=x^2 , v = cos nx

OpenStudy (anonymous):

uv-∫vdu u mean this?

hartnn (hartnn):

yes. u = x^2 what will u take dv =... ?

OpenStudy (anonymous):

actually dv=cosnx and v=-sinnx

hartnn (hartnn):

yes, yes, go ahead and solve it...

hartnn (hartnn):

you'll again need product rule when u solve for sin nx (2x)dx

OpenStudy (anonymous):

yea thnx... i'll try and finally put the answer here to prove;)

hartnn (hartnn):

ok.